OpenStudy (anonymous):

Question about derivative of ln(x*sqrt(x^2-6) Would it be derivative of lnx + lnsqrt(x^2-6) which would be 1/x + (1/2) (1/sqrtx^2-6) (2x)

OpenStudy (anonymous):

i think so

OpenStudy (anonymous):

\[\ln(x\sqrt{x^2-6})=\ln(x)+\frac{1}{2}\ln(x^2-6)\]

OpenStudy (anonymous):

taking derivatives gives \[\frac{1}{x}+\frac{x}{x^2-6}\]

OpenStudy (anonymous):

I'm getting confused with x/x^2-6

OpenStudy (anonymous):

wouldn't it be 1/sqrtx^2-6